# Minimal Algorithmic Information Loss Methods for Dimension Reduction, Feature Selection and Network Sparsification

### v11

@@ Abstract
We present a novel, domain-agnostic, model-independent, unsupervised, and universally applicable approach for data summarization. Specifically, we focus on addressing the challenge of reducing certain dimensionality aspects, such as the number of edges in a network, while retaining essential features of interest. These features include preserving crucial network properties like degree distribution, clustering coefficient, edge betweenness, and degree and eigenvector centralities. Our approach outperforms state-of-the-art network reduction techniques by achieving an average improvement in feature preservation. Previous methods grounded in statistics or classical information theory have been limited in their ability to capture more intricate patterns and features, particularly nonlinear patterns stemming from deterministic computable processes. Moreover, these approaches heavily rely on a priori feature selection, demanding constant supervision. Our findings demonstrate the effectiveness of the algorithms proposed in this study in overcoming these limitations, all while maintaining a time-efficient computational profile. In many instances, our approach not only matches but also surpasses the performance of established network reduction algorithms. Furthermore, we extend the applicability of our method to lossy compression tasks involving images or any bi-dimensional data. This highlights the versatility and broad utility of our approach in various domains.
Keywords: data summarisation; lossy compression; data dimensionality reduction; network complexity; image segmentation; algorithmic machine learning
@@ 1 Introduction
The study of large and complex datasets, or Big Data, organised as networks, has emerged as one of the central challenges in most areas of science and technology, cellular and molecular networks in biology being one of the prime examples. The analysis of large-scale networks poses significant computational and analytical challenges. Hence, network sparsification methods have emerged as indispensable tools to address these challenges by reducing the complexity of networks while preserving essential structural or functional properties [14, 19, 21, 27, 31, 35, 46]. Moreover, it’s worth noting that data reduction techniques extend beyond network data. Methods for image sparsification also play pivotal roles in managing and analysing large datasets. In essence, data reduction, in a broader sense, involves transforming various types of numerical, textual, or visual digital information into more compact representations while retaining specific properties of significance.
@@
This overarching concept of data reduction encompasses a wide range of methods, including network sparsification, image compression, and simplifying cellular automata rules. The common thread among these methods is the pursuit of more efficient data representation, enabling streamlined analysis and storage across diverse fields. The primary objective of data dimensionality reduction is twofold, much like two sides of the same coin. One facet emphasises minimising information loss, ensuring that valuable insights and nuances are preserved as data is transformed into a more manageable form. The other aspect underscores the importance of preserving the most ’meaningful’ features that uniquely characterise an object, a concept commonly referred to as feature selection. This dual objective serves as a fundamental principle in data reduction, often constituting a trade-off to be tackled in the data analysis phase. Traditionally, identifying these meaningful features relies on an user-centric, subjective criterion. Specific analytical goals and human expertise guide selection processes. For example, linear algebraic (e.g., matrix analysis) and statistically-based dimensionality reduction techniques attempt to minimise statistical information loss under certain algebraic (interpreted as signal and noise) conditions, consequently maximizing the statistical mutual information between the desired information and the dimensionally-reduced output.
@@
The primary objective of dimensionality reduction in network analysis is to approximate a complex network with a sparser representation. Numerous methods for graph sparsification, a form of graph summarization, have been documented in the literature [46]. Some earlier approaches, such as Chew’s [27], relied on criteria like the shortest path distance between pairs of vertices for network sparsification. Others, like Benczur and Karger [20], utilized cut problems as a basis for sparsification. More recently, techniques such as spectral similarity of graph Laplacians have been employed by Spielman et al. [58]. When it comes to network dimensionality reduction, one must select criteria for preserving graph-theoretic properties like graph distance, clustering coefficient, or degree distribution, or combinations thereof. However, it’s important to acknowledge that no finitely computable approach can comprehensively identify all possible features of interest in a dataset. This includes recursively enumerable features that the set of all Turing machines can characterise concurrently [65]. Consequently, data analysts often face the daunting task of making arbitrary choices regarding which features to prioritize, as exemplified in previous research [66, 67].
@@
This paper’s main goal is to develop an unsupervised method that remains theoretically optimal and is independent of computational models, programming languages, physical implementations, and available resources. Furthermore, this method should be agnostic to selecting features of interest, encompassing all computably enumerable parts within a given dataset earmarked for reduction. We introduce a family of semi-computable algorithms designed to preserve computable properties, encompassing both statistical and algorithmic aspects. This approach is a generalization of various dimension reduction techniques since statistically definable, generable, or recognizable patterns constitute a subset of those that algorithmic approaches can theoretically handle.
@@
Our method represents a novel approach to designing theoretically optimal lossy compression techniques, drawing inspiration from principles and estimations related to theoretical optimal lossless compression [63, 18]. By employing efficient (polynomial) approximations to algorithmic complexity values from recent numerical methods of algorithmic probability [32, 56, 68, 55, 62, 69], we demonstrate how these algorithms can preserve structure properties, performing similarly to, if not consistently outperforming, state-of-the-art algorithms in e.g. the area of network dimension reduction. We test our algorithms on non-trivial cases against transitive and non-linear (spectral) methods involving simple graphs where statistical regularities are even easier to conceal more intricate computable processes or embedded encodings, and thus may easily fool weaker, linear and computable measures [66, 65, 11, 12].
@@
We will demonstrate how our method preserves critical network topological properties, including degree distribution, clustering coefficient, edge betweenness, degree, and eigenvector centralities compared to other methods. These fundamental properties provide essential insights into network structure and functionality. Degree distribution, for instance, reveals how nodes are interconnected, helping us understand the overall network architecture and the distribution of connections. It is indispensable for identifying hubs and assessing network robustness. The clustering coefficient quantifies the tendency of nodes to form tightly connected clusters. This metric is invaluable for detecting network communities and assessing resilience against failures or attacks. Edge betweenness measures the significance of edges in facilitating efficient communication between nodes. Identifying bottleneck edges is vital for understanding network flow and vulnerability. Degree and eigenvector centralities help pinpoint influential nodes within a network. High-degree nodes are hubs, while high-eigenvector centrality nodes are well-connected to other influential nodes. These centralities are vital for identifying key players in various network types.
@@
By preserving these critical properties, our approach strives to provide more meaningful and accurate representations of complex networks [67, 11, 12] while addressing the challenges inherent in data reduction techniques. Central to our approach are perturbations in the form of edge deletions aimed at minimizing the loss of algorithmic information content [13]. We have aptly named our algorithm MILS, signifying ’Minimal Information Loss Selection,’ as it encapsulates our mission of preserving essential information while reducing data dimensionality. Notably, the algorithms presented here remain agnostic to the specific method for approximating algorithmic complexity. They can be implemented using various approaches, including entropy or lossless compression. However, our description employs a technique based on [68, 72, 63], which not only encompasses but also enhances other statistics- and entropic-based methods due to its intrinsic structural characteristics.
@@
In Section 2.2, where we detail the algorithm, it’s essential to recognise that substituting the underlying methods for approximating algorithmic information content with alternatives like Shannon entropy or lossless compression algorithms represents special cases within the broader algorithm based on algorithmic complexity. This comprehensive approach effectively covers all these less powerful cases because the block decomposition method (BDM) calculation we employ subsumes entropy-based approximations [68]. BDM presents a novel approach to understanding and quantifying the complexity of biological networks and systems. BDM extends the power of the Coding Theorem Method (CTM) [62, 55], offering a closer connection to algorithmic complexity than previous methods based on statistical regularities. This method effectively decomposes complex objects into smaller components using a set of short computer programs. These components are then sequenced to reconstruct the original object, allowing for efficient estimations of algorithmic complexity. BDM as has been shown by our team and other researcher addresses a fundamental challenge in science: capturing the complexity of objects for classification and profiling. It bridges the gap between computable measures like Shannon entropy [55], which lacks invariance to object descriptions and uncomputable universal estimates of complexity. The method lies between these two, dividing data into smaller pieces to partially circumvent the uncomputability problem. This results in efficient estimations, albeit with some loss of accuracy.
@@
BDM has been applied in various fields and applications outperforming other complexity measures [28, 67], for example to evolutionary systems [38], artificial intelligence [41, 17], and the detection of bio- and technosignature [26, 29]. Significantly, BDM’s application to graphs and causal calculus has enabled the deconvolution of complex network behaviors, identifying distinct topological generating mechanisms [33, 63, 18, 61]. This approach uses a perturbation-based causal calculus to infer model representations [13, 60, 40], effectively tackling the challenges of causation in complex systems. It has been demonstrated by us and others how by analyzing data from discrete dynamical systems like cellular automata and complex networks, BDM-based algorithms provides a deeper understanding of the intricate interplay of elements within these systems, revealing hidden patterns and relationships, and contributing significantly to the field of algorithmic complexity in biological networks[33, 18, 64, 70, 72, 63]. Moreover, algorithmic complexity, being an asymptotically optimal measure of irreducible information content [26, 34, 43, 24], holds the unique advantage of capturing algorithmically determined patterns that defy reduction to statistical patterns [28, 66].
@@
Our implementation of the MILS algorithm, rooted in algorithmic complexity, offers the added benefit of avoiding certain low-complexity deceptive phenomena. Traditional methods based solely on entropy and statistics are prone to such issues [28, 66]. Significantly, this avoidance does not compromise running time efficiency or data size reduction performance, as evidenced by our results in Section 4. These results demonstrate that we either match or, more often than not, surpass the performance of the best current algorithms, both for local and global graph properties.
@@
The approach presented in this article lies in its potential to assess the effectiveness of all reduction techniques while striving for optimal reduction by minimising algorithmic information loss. This represents a nonlinear generalisation encompassing a broader range of techniques beyond merely preserving statistical or domain-specific algebraic properties. While our focus in this article centers on edge deletions in graphs to describe the MILS algorithm, it’s essential to emphasise that MILS can be applied to a wide array of objects and various types of constructive or destructive perturbations [13, 60, 72, 70] on their constituent elements. For instance, MILS can seamlessly extend its applicability to node deletion, offering versatility in addressing diverse algorithmic perturbations [13] aimed at minimising information differences between objects.

### v12

@@ Abstract
We present a novel, domain-agnostic, model-independent, unsupervised, and universally applicable Machine Learning approach for dimensionality reduction based on the principles of algorithmic complexity. Specifically, but without loss of generality, we focus on addressing the challenge of reducing certain dimensionality aspects, such as the number of edges in a network, while retaining essential features of interest. These features include preserving crucial network properties like degree distribution, clustering coefficient, edge betweenness, and degree and eigenvector centralities but can also go beyond edges to nodes and weights for network pruning and trimming. Our approach outperforms classical statistical Machine Learning techniques and state-of-the-art dimensionality reduction algorithms by preserving a greater number of data features that statistical algorithms would miss, particularly nonlinear patterns stemming from deterministic recursive processes that may look statistically random but are not. Moreover, previous approaches heavily rely on a priori feature selection, which requires constant supervision. Our findings demonstrate the effectiveness of the algorithms in overcoming some of these limitations while maintaining a time-efficient computational profile. Our approach not only matches, but also exceeds, the performance of established and state-of-the-art dimensionality reduction algorithms. We extend the applicability of our method to lossy compression tasks involving images and any multi-dimensional data. This highlights the versatility and broad utility of the approach in multiple domains.
keywords:
Machine Learning , lossy algorithmic complexity, recursive compression, data dimensionality reduction, network complexity, algorithmic image segmentation
journal: Information Sciences
[…]
[KCL]organization=Department of Biomedical Computing, Deparment of Digital Twins, School of Biomedical Engineering and Imaging Sciences, addressline=King’s Institute for AI, King’s College London, country=U.K.
[…]
[Turing]organization=The Alan Turing Institute, addressline=British Library, country=U.K.
[…]
[OxfImAl]organization=Oxford Immune Algorithmics, Oxford University Innovation and London Institute for Healthcare Engineering, country=U.K.
[…]
[ADL]organization=Algorithmic Dynamics Lab and Department of Oncology-Pathology, Center of Molecular Medicine, addressline=Karolinska Institutet, country=Sweden
[…]
[Cross]organization=Cross Labs, addressline=Cross Compass, country=Japan
[…]
[OncPat]organization=Karolinska Institutet, addressline=Department of Oncology-Pathology, Center of Molecular Medicine, city=Stockholm, country=Sweden
[…]
[WisMad]organization=University of Wisconsin-Madison, addressline=John W. and Jeanne M. Rowe Center for Research in Virology, Morgridge Institute for Research, country=USA
[…]
[Unicamp]organization=University of Campinas (UNICAMP), addressline=Centre for Logic, Epistemology and the History of Science (CLE), country=Brazil
[…]
[LNCC]organization=National Laboratory for Scientific Computing (LNCC), addressline=Data Extreme Lab, country=Brazil
[…]
[Vene1]organization=Universidad Central de Venezuela, addressline=Instituto Nacional de Bioingeniería, country=Venezuela
[…]
[Vene2]organization=Universidad Central de Venezuela, addressline=Department of Mathematics, country=Venezuela
[…]
[KAUST]organization=King Abdullah University of Science and Technology (KAUST), addressline=Biological and Environmental, Science and Engineering Division, country=Kingdom of Saudi Arabia
@@ 1 Introduction
The study of large and complex datasets such as networks, has emerged as one of the central challenges in most areas of complex systems and Machine Learning, cellular and molecular networks in biology being one of the prime examples. The analysis of large-scale networks poses significant computational and analytical challenges.
@@
Network sparsification methods have become indispensable tools[1, 2] for addressing these challenges by reducing the complexity of networks[3, 4] while preserving essential structural [5] or functional properties [6, 7]. Data reduction techniques extend beyond network data. Methods for image sparsification also play an essential role in managing and analysing large datasets. In essence, data reduction, in a broader sense, involves transforming various types of numerical, textual, or visual digital information into more compact representations while retaining specific properties of significance.
This overarching concept of data reduction encompasses a wide range of methods, including network sparsification, image compression, and simplifying cellular automata rules. The common thread among these methods is the pursuit of more efficient data representation, enabling streamlined analysis and storage across various fields.
@@
The primary objective of data dimensionality reduction in Machine Learning is twofold, much like two sides of the same coin. One facet emphasizes minimizing information loss, ensuring that valuable insights and nuances are preserved as data is transformed into a more manageable form. The other aspect underscores the importance of preserving the most ‘meaningful’ features that uniquely characterise an object, a concept commonly referred to as feature selection. This dual objective serves as a fundamental principle in data reduction, often constituting a trade-off to be tackled in the data analysis phase. Traditionally, identifying these meaningful features is based on a subjective criterion that is user-centric. Specific analytical goals and human experience guide the selection processes. For example, linear algebraic (e.g., matrix analysis) and statistically-based dimensionality reduction techniques attempt to minimise statistical information loss under certain algebraic (interpreted as signal and noise) conditions, consequently maximizing the statistical mutual information between the desired information and the dimensionally-reduced output.
@@
The primary objective of dimensionality reduction in network analysis is to approximate a complex network with a sparser representation. Numerous methods for graph sparsification, a form of graph summarisation, have been documented in the literature [7]. Some earlier approaches, such as Chew’s [4], relied on criteria like the shortest path distance between pairs of vertices for network sparsification. Others, like Benczur and Karger [8], utilized cut problems as a basis for sparsification. Techniques such as spectral similarity of graph Laplacians have been employed by Spielman et al. [9]. When it comes to network dimensionality reduction, one must select criteria for preserving graph-theoretic properties like graph distance, clustering coefficient, or degree distribution, or combinations thereof. However, it’s important to acknowledge that no finitely computable approach can comprehensively identify all possible features of interest in a dataset. This includes recursively enumerable features that the set of all Turing machines can characterise concurrently [10]. Consequently, data analysts often face the daunting task of making arbitrary choices regarding which features to prioritize, as exemplified in previous research [11, 12].
This paper’s main goal is to develop an unsupervised method that is theoretically sound and optimal in distinguishing noise from data, efficient, and independent from traditional statistical Machine Learning approaches based on classical information theory. Furthermore, the methods introduced are agnostic in selecting features of interest, encompassing all estimated computably enumerable parts within a given dataset earmarked for reduction.
@@
We introduce a family of semi-computable algorithms designed to preserve computable properties, encompassing both statistical and algorithmic aspects. This approach is a generalisation of various dimension reduction techniques since statistically definable, generable, or recognizable patterns constitute a subset of those that algorithmic approaches can theoretically handle.
@@
Our method represents a novel approach to designing theoretically optimal lossy compression techniques, drawing inspiration from principles and estimations related to theoretical optimal lossless compression [14, 13]. By employing efficient (polynomial) approximations [16, 15] to algorithmic complexity [17] values from recent numerical methods of algorithmic probability, like in the works of  [18] [19] and [20], we demonstrate how these algorithms can preserve structural properties, performing similarly to, if not consistently outperforming, state-of-the-art algorithms in, e.g., the area of network dimension reduction. We test our algorithms on non-trivial cases against transitive and non-linear (spectral) methods involving simple graphs where statistical regularities are even easier to conceal more intricate computable processes or embedded encodings, and thus may easily fool weaker [11], linear and computable measures [10, 21].
@@
We will demonstrate how our method preserves critical network topological properties, including degree distribution, clustering coefficient, edge betweenness, degree, and eigenvector centralities compared to other methods. These fundamental properties provide essential information about the structure and functionality of the network. Degree distribution, for instance, reveals how nodes are interconnected, helping us understand the overall network architecture and the distribution of connections. It is indispensable for identifying hubs and assessing network robustness. The clustering coefficient quantifies the tendency of the nodes to form tightly connected clusters. This metric is invaluable for detecting network communities and assessing resilience against failures or attacks. Edge betweenness measures the significance of edges in facilitating efficient communication between nodes. Identifying bottleneck edges is vital to understanding network flow and vulnerability. The degree and eigenvector centralities help identify influential nodes within a network. High-degree nodes are hubs, while high-eigenvector centrality nodes are well-connected to other influential nodes. These centralities are vital for identifying key players in various types of networks.
@@
By preserving these critical properties, our approach strives to provide more meaningful and accurate representations of complex networks [12, 22] while addressing the challenges inherent in data reduction techniques. Central to our approach are perturbations in the form of edge deletions aimed at minimizing the loss of algorithmic information content [23]. We have aptly named our algorithm MILS, signifying ’Minimal Information Loss Selection,’ as it encapsulates our mission of preserving essential information while reducing data dimensionality. Notably, the algorithms presented here remain agnostic to the specific method for approximating algorithmic complexity. They can be implemented using various approaches, including entropy or lossless compression. However, our description employs a technique based on [24, 14], which not only encompasses but also enhances other statistics- and entropic-based methods due to its intrinsic structural characteristics.
@@
It is essential to notice that substituting the underlying methods to approximate algorithmic information content with alternatives such as Shannon entropy or lossless compression algorithms represents special cases within the broader algorithm based on algorithmic complexity. This comprehensive approach effectively covers all these less powerful cases, because the block decomposition method (BDM) calculation we employ subsumes entropy-based approximations [16]. BDM presents a novel approach to understanding and quantifying the complexity of biological networks and systems. BDM extends the power of the Coding Theorem Method (CTM) [20, 15], offering a closer connection to algorithmic complexity than previous methods based on statistical regularities. This method effectively decomposes complex objects into smaller components using a set of short computer programs. These components are then sequenced to reconstruct the original object, allowing for efficient estimations of algorithmic complexity. BDM as has been shown by our team and other researcher addresses a fundamental challenge in science: capturing the complexity of objects for classification and profiling. It bridges the gap between computable measures like Shannon entropy [15], which lacks invariance to object descriptions and uncomputable universal estimates of complexity. The method lies between these two, dividing data into smaller pieces to partially circumvent the uncomputability problem. This results in efficient estimations, albeit with some loss of accuracy.
@@
BDM has been applied in various fields and applications that outperform other complexity measures [25, 12], for example, to evolutionary systems [26], artificial intelligence [28, 27], and the detection of bio- and technosignature [29]. Significantly, the application of BDM to graphs and causal calculus has enabled the deconvolution of complex network behaviors [30], identifying distinct topological generating mechanisms [13]. This approach uses a perturbation-based causal calculus to infer model representations [31, 32], effectively tackling the challenges of causation in complex systems. We and others have demonstrated how by analysing data from discrete dynamical systems such as cellular automata and complex networks, BDM-based algorithms provide a deeper understanding of the intricate interplay of elements within these systems, revealing hidden patterns and relationships, and contributing significantly to the field of algorithmic complexity [33] in biological networks [34, 24]. Moreover, algorithmic complexity, being an asymptotically optimal measure of irreducible information content [35, 36], has the unique advantage of capturing algorithmically determined patterns that defy reduction to statistical patterns [25].
@@
Our implementation of the MILS algorithm, rooted in algorithmic complexity, offers the added benefit of avoiding certain low-complexity deceptive phenomena. Traditional methods based solely on entropy and statistics are prone to such issues [25, 11]. Significantly, this avoidance does not compromise running time efficiency or data size reduction performance for graphs, as evidenced by our results in Section 5. On the other hand, when implemented in ML pipelines, computational time still presents challenges. These results demonstrate that we either match or, more often than not, surpass the performance of the best current algorithms, both for local and global graph properties. In image-compression tasks, we show that MILS presents the highest mean classification accuracy per bit of compressed images when such images are used in a particular classification task.
@@
The approach presented in this article lies in its potential to assess the effectiveness of all reduction techniques while striving for optimal reduction by minimizing algorithmic information loss. This represents a nonlinear generalisation encompassing a broader range of techniques beyond merely preserving statistical or domain-specific algebraic properties. While our focus in this article centers on edge deletions in graphs to describe the MILS algorithm, it’s essential to emphasise that MILS can be applied to a wide array of objects and various types of constructive or destructive perturbations [24, 34] on their constituent elements. For instance, MILS can seamlessly extend its applicability to node deletion, offering versatility in addressing diverse algorithmic perturbations [23] aimed at minimizing information differences between objects.
2 Preliminaries on complexity measures and algorithmic probability
A comprehensive discussion of complexity measures and algorithmic probability is provided in the Supplementary Material. Meanwhile, this section introduces key preliminaries to establish the foundational concepts necessary for the main discussion. Using the principles of classical information theory, Block Decomposition Method (BDM) or BDM combines the calculation of the global Shannon Entropy rate of the object with local estimations to algorithmic complexity of smaller blocks in which the object is decomposed for which values are found in a pre-computed database of direct approximations of algorithmic probability [16].
The BDM relies on the following assumptions:
  1. 1.
    In the case of small enough objects (e.g., binary strings), their algorithmic complexity can be approximated using exhaustive search.
  2. 2.
    For larger objects, breaking them into smaller parts allows for the approximation of the overall complexity by summing the complexity of individual blocks, with a correction factor to account for interactions between the blocks.
  3. 3.
    For every other length, values of Shannon Entropy rates are calculated and combined with the previous values by using the same principles of information theory.
Formally, let $x$ be a string divided into blocks $x_{i}$, with $x=x_{1}\oplus x_{2}\oplus\dots\oplus x_{n}$, where $\oplus$ denotes a concatenation operator. The BDM complexity of $x$, denoted $\text{BDM}(x)$, is:
$\text{BDM}(x)=\sum_{i=1}^{n}\text{CTM}(x_{i})+\log n$
Where:
  • $\text{CTM}(x_{i})$ is the algorithmic complexity approximation for block $x_{i}$, derived from the Coding Theorem Method (CTM).
  • $\log n$ is a correction factor accounting for the interactions between the blocks.
The Coding Theorem Method (CTM) is a method based on the Coding Theorem and Algorithmic Probability [36], which connects classical probability to Kolmogorov complexity. The CTM maps sets of micro programs (e.g., small Turing machines) to small assembly objects for which it can empirically estimate the algorithmic probability of an object based on the following relationship [15]:
$K(s)\approx-\log P(s)$
Where:
  • $K(s)$ is the Kolmogorov complexity of string $s$.
  • $P(s)$ is the algorithmic probability of string $s$, as defined by Solomonoff’s distribution.
The CTM leverages pre-computed distributions from small Turing machines to approximate the algorithmic complexity of small strings by estimating their algorithmic probability [19, 20]. BDM and CTM can be seen as practical approximations of algorithmic probability and Kolmogorov complexity, especially in scenarios where exact computation is impossible. The relationship can be summarised as follows:
On the one hand, CTM provides an approximation to algorithmic probability $P(s)$ by connecting the empirical frequency of a string’s generation in a small Turing machine model to its Kolmogorov complexity $K(s)$, using the relation $K(s)\approx-\log P(s)$. On the other hand, BDM offers a method to map micro programs to their respective small pieces of an object requiring an algorithmic explanation.
2.1 Algorithmic information theory, data compression and feature selection
BDM, CTM and Algorithmic Probability offer a principled approach to data compression and feature selection by quantifying algorithmic complexity rather than relying on traditional statistical methods. In data compression, these methods allow for data decomposition into patches and estimating their Kolmogorov complexity, ultimately selectively removing redundant regions while preserving structurally meaningful information. This empowers compression routines which prioritize algorithmically significant structures rather than merely frequent ones, resulting in compression that minimises storage requirements while retaining essential data properties.
In feature selection, these methods provide a more fundamental measure of information content than conventional statistical metrics like correlation or variance. These methods help identify which features contain essential computational information, enabling the removal of redundant or low-complexity features without significant loss of structure. Thus, their ability to optimize data representation while retaining essential information makes them valuable tools for scientific computing, machine learning, and complex data-driven applications.